OpenStudy (anonymous):

Find the general solution of the equation Sec^2 (2x) = 1 - tan2x

OpenStudy (anonymous):

sec^2(2x)=1+tg^2(2x) 1+tg^2(2x)=1-tg(2x) tg^2(2x)+tg(2x)=0 tg(2x)[tg(2x)+1]=0 tg(2x)=0 or tg(2x)+1=0 2x=arctg(0) or 2x=arctg(-1)

OpenStudy (anonymous):

\[ \sec^2 (2x) = 1 - \tan2x\\ 1 + \tan^2(2x)=1 -\tan(2x)\\ \tan^2(2x)+\tan(2x)=0\\ \tan(2x)( \tan(2x) +1)=0\\ \tan(2x) =0\\ \tan(2 x) =-1 \]

OpenStudy (anonymous):

\[ \tan(2x)=0\\ 2x =k \pi\\ x= \frac {k \pi}2 \]

OpenStudy (anonymous):

\[ \tan(2 x) =-1\\ \tan(2 x) = \tan \left ( - \frac \pi 4 \right)\\ x = -\frac \pi 8 + k \pi \]